OpenStudy (anonymous):

A certain cell culture grows at a rate proportional to the number of cells present. If the culture contains 500 cells initially and 800 after 24 hours, how many cells will be present after a further 12 hours?

OpenStudy (anonymous):

let x = number of cells at time t dx/dt = kx solving this differential equation gives x =500 e^(kt) when x = 800, t = 1 (day): 800 = 500e^k k = 0.47 so x = 500e^(0.47t) so after another 12 hours when t = 1.5 x = 500e^(0.47*1.5) x = 1012

OpenStudy (anonymous):

1012